OpenStudy (anonymous):

Evaluate -5x^2 − 8x + 1 when x = -3

OpenStudy (anonymous):

-25

OpenStudy (anonymous):

Answer=-20. Solution is as follows: -5(-3)(-3)-8(-3)+1=-5(9)+25=-45+25=-20

OpenStudy (anonymous):

yea sorry.. made a mistake

OpenStudy (anonymous):

all good